Math 246, Section 1301

Differential Equations for Scientists and Engineers

Fall, 2001

Prerequisites

Textbooks

Lectures

Instructor

Grader

Homework

Exams

Grading Math 141 (Calculus II) or equivalent.

1. Elementary Differential Equations, Seventh edition, by W. E. Boyce and R. C. DiPrima, John Wiley & Sons, Inc., 2001.

2. Differential Equations with Matlab, by Kevin R. Coombes, B. R. Hunt, R. L. Lipsman, J. E. Osborn, and G. J. Stuck, John Wiley & Sons, Inc., 2000.

MWF 2:00–2:50, Math 0103.

Bo Li (Office hours: W 3:00–4:00, Th 11:00–12:00; Office: Math 4304; Phone: 301-405-5088; E-mail: bli@math.umd.edu).

Sunhee Kim (E-mail: sunny@math.umd.edu).

Assigned regularly and graded partially. There are two types of homework: one is analytical and the other Matlab related. For an analytical type of homework, exercise problems are chosen from the first textbook Elementary Differential Equations. For a Matlab related homework, exercise problems are chosen from the second textbook Differential Equations with Matlab. Note: no late homework will be accepted unless a written verification of a valid excuse (such as hospitalization, family emergency, religious observance, court appearance, etc.) is provided and approved by the instructor.

There will be two midterm and one final, close-book and close-note exams. The final exam will be cumulative.

Midterm exam 1: 2:00–2:50, Friday, October 5, Math 0103.

Midterm exam 2: 2:00–2:50, Friday, November 16, Math 0103.

Final exam: 1:30–3:30, Thursday, December 13. Note: neither rescheduled nor make-up exams will be allowed unless a written verification of a valid excuse (such as hospitalization, family emergency, religious observance, court appearance, etc.) is provided and approved by the instructor.

The final course grade will be determined based on the exams and homework with the following weights: final exam — 40%, midterm exam 1 — 20%, midterm exam 2 — 20%, homework of analytical type — 10%, homework related to Matlab — 10%. The final grade will be curved.